OpenStudy (anonymous):

Use the following information to solve problems 1 and 2. One-half of the students at Greendale Middle School take art. Two-thirds of the students take Spanish. Shawn uses a coin and a standard number cube as a simulation of the students at Greendale Middle School. Shawn letts "T" represent a student taking art and "H" represent a student not taking art. He lets the numbers 1, 2, 3, and 4 on the number cube represent a student taking Spanish, and the numbers 5 and 6 represent a student not taking Spanish. The table shows the results of Deshawn's samulation

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

Find the probability that a student at Greendale Middle School takes both art and Spanish

OpenStudy (triciaal):

I saw this question yesterday, sorry I don't understand the table

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

@paki

OpenStudy (triciaal):

probability of taking art = 1/2 probability of taking spanish = 4/6 = 2/3

OpenStudy (anonymous):

a 3/20 b 2/5 c 7/20 d 1/10

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@triciaal do u know the answer

OpenStudy (triciaal):

not yet finally just figured out the table

OpenStudy (anonymous):

oh ok

OpenStudy (triciaal):

20 students in the class

OpenStudy (anonymous):

yup

OpenStudy (triciaal):

example T5 means takes art but not spanish etc

OpenStudy (anonymous):

yes

OpenStudy (triciaal):

suggest fill in the venn diagram to find the number of students in each section then go from there

OpenStudy (anonymous):

@asnaseer

OpenStudy (triciaal):